OpenStudy (anonymous):

(6, 3) and (8, 6). Part 1: Slope (m) = change in y over change in x. Part 2: Now find the equation using the point slope formula: Point Slope Formula = y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2 Part 4: Give an example of an equation for a line perpendicular to your answer from Part 2

OpenStudy (anonymous):

Part 1: m= 6 - 3/8 - 6 = 3/2 = 1.5 3/2 = m (Slope) Part 2: Point Slope Formula = y – y1 = m (x – x1) y - 3 = 3/2 (x - 6) y - 3 = x + 6 y- 3 +3 = x +6 +3 y = x + 9